# GaussianFocus: Constrained Attention Focus for 3D Gaussian Splatting

### v1

@@ Abstract
Recent developments in 3D reconstruction and neural rendering have significantly propelled the capabilities of photo-realistic 3D scene rendering across various academic and industrial fields. The 3D Gaussian Splatting technique, alongside its derivatives, integrates the advantages of primitive-based and volumetric representations to deliver top-tier rendering quality and efficiency. Despite these advancements, the method tends to generate excessive redundant noisy Gaussians overfitted to every training view, which degrades the rendering quality. Additionally, while 3D Gaussian Splatting excels in small-scale and object-centric scenes, its application to larger scenes is hindered by constraints such as limited video memory, excessive optimization duration, and variable appearance across views. To address these challenges, we introduce GaussianFocus, an innovative approach that incorporates a patch attention algorithm to refine rendering quality and implements a Gaussian constraints strategy to minimize redundancy. Moreover, we propose a subdivision reconstruction strategy for large-scale scenes, dividing them into smaller, manageable blocks for individual training. Our results indicate that GaussianFocus significantly reduces unnecessary Gaussians and enhances rendering quality, surpassing existing State-of-The-Art (SoTA) methods. Furthermore, we demonstrate the capability of our approach to effectively manage and render large scenes, such as urban environments, whilst maintaining high fidelity in the visual output. (The link to the code will be made available after publication)
keywords:
3D Reconstruction , 3D Gaussian Splatting , Neural Rendering ,  Novel View Synthesis
journal: Knowledge-Based Systems
[…]
[1] organization=School of Electrical and Data Engineering, University of Technology Sydney, addressline=15 Broadway, city=Ultimo, postcode=2007, state=NSW, country=Australia
@@ 1 Introduction
Novel View Synthesis (NVS) is fundamental for modern computer graphics and vision, extending to virtual reality, autonomous driving, and robotics. Primitive-based models such as meshes and point clouds [1, 2, 3], optimized for GPU rasterization, deliver fast but often lower-quality images with discontinuities. The introduction of Neural Radiance Fields (NeRF) by [4] marked a significant advancement, employing a multi-layer perceptron (MLP) to achieve high-quality, geometrically consistent renderings of new viewpoints. However, NeRF’s reliance on time-consuming stochastic sampling can lead to slower performance and potential noise issues.
Recent advancements in 3D Gaussian Splatting (3DGS) [5] have significantly enhanced rendering quality and speed. This technique refines a set of 3D Gaussians, initialized using Structure from Motion (SfM) [6], to model scenes with inherent volumetric continuity. This facilitates fast rasterization by projecting them onto 2D planes. However, 3DGS often produces artifacts when camera viewpoints deviate from the training set and lack detail during zooming. To address these issues, newer models [7, 8] employ a 3D smoothing filter to regularize the frequency distribution and utilize anchor points to initialize 3D Gaussians, thereby enhancing visual accuracy and applicability in diverse scenarios. Despite these advances, 3DGS-based models still tend to produce oversized Gaussian spheres that ignore scene structure, leading to redundancy and scalability issues in complex environments. Additionally, these models struggle with detail reconstruction, particularly at edges and high-frequency areas. This often leads to suboptimal rendering quality. Moreover, reconstructing large-scale scenes like towns or cities represents a significant challenge due to GPU memory constraints and computational demands. To mitigate these problems, models often reduce training input randomly, which compromises reconstruction quality and results in incomplete outcomes.
@@
    We propose a 3DGS-based patch attention algorithm with novel edge and frequency losses designed to enhance the details and reduce spatial frequency artifacts within scene reconstructions. This improves the detailing quality and intricacy of the rendered scenes.
@@
    We impose constraints on overly large Gaussian spheres to mitigate the occurrence of “air walls”, thus refining the scene reconstruction’s fidelity and enhancing the granularity of the resulting models. Moreover, these constraints allow the achievement of superior reconstruction results with fewer training iterations.
@@
    For large-scale scene reconstruction, our approach involves subdividing the scene for subdivision-based reconstruction and subsequent recombination. This method addresses the challenge posed by existing 3DGS-based models that fail to directly reconstruct extensive scenes, thereby enhancing the scalability and applicability of our reconstruction framework.
In this paper, we structure the content as follows: Section 3 outlines the methods we employed. In Section 4, we present our experimental framework compare its performance to other advanced 3DGS-based models and discuss the ablation studies, implementation details and model limitations. We conclude the paper in Section 5.
@@ 2 Related Work
Volumetric Rendering methods    Volumetric approaches utilize structures such as multiplane images, voxel grids or neural network models to depict scenes as continuous functions that define their volume, density, and colour characteristics. The introduction of Neural Radiance Fields (NeRF) [4] marked a significant advancement in scene representation technology. This method employs a multilayer perceptron (MLP) to parameterize a continuous volumetric function. This parameterization facilitates the creation of photorealistic images that exhibit precise details. These details and effects are dependent on the viewer’s perspective, achieved through volumetric ray tracing. Nevertheless, the application of the vanilla NeRF model is hindered by its high demand for computational power and memory. To overcome these challenges, subsequent research has sought to refine NeRF’s efficiency and extend its scalability. Such improvements have been achieved through the implementation of discretized or sparse volumetric frameworks, such as voxel grids and hash tables. These frameworks [10, 11, 12, 13, 14, 15] are crucial as they hold learnable features that act as positional encodings for 3D coordinates. Additionally, these methods employ hierarchical sampling techniques [16, 17, 18] and utilize low-rank approximations [10]. Despite these enhancements, the dependence on volumetric ray marching continues, which leads to compatibility challenges with traditional graphics equipment and systems primarily engineered for polygonal rendering. Additionally, recent innovations have adjusted NeRF’s approach to geometry and light emission representation, improving the rendering of reflective surfaces [19] and enabling more effective scene relighting by separately addressing material and lighting attributes [20, 21, 22]. Moreover, certain approaches implement rendering methods that include techniques like interpolating between different viewpoints [23], or they employ depth sensing technology to accurately acquire the geometry of a scene [24]. Nonetheless, these methodologies are generally confined to indoor settings. This limitation arises from the constraints on the sensor’s operational range and its inherent design, which is primarily geared towards manipulating 2D scenes.
Point-based Rendering methods    Point-based rendering methods leverage point clouds as fundamental geometric units for the visualization of scenes. The typical methods [25, 26] involve using graphical APIs and GPU-specific modules to rasterize these unstructured point sets at a constant size. Despite the rapid rendering and flexibility in managing changes in topology, this method is prone to forming holes and outliers, which frequently result in rendering artifacts. To address these gaps, research on differentiable point-based rendering has become prevalent, aiming to precisely model the geometry of objects [27, 28, 3, 29, 30]. Research has examined the use of differentiable surface splatting in studies like [3, 30], in which points are interpreted as larger-than-one-pixel geometric objects such as surfels, elliptic shapes, or spheres. Methods [31, 32] have enriched point features with neural network capabilities and processed them through 2D CNNs for visualization. In contrast, Point-NeRF [33] has demonstrated superior capabilities in synthesizing new views of high quality using volume rendering, incorporating strategies like region growth and point reduction during its optimization phase. However, this technique is limited by its dependence on volumetric ray-marching, impacting its display speed. Remarkably, the 3DGS [5] framework employs directionally dependent 3D Gaussians for three-dimensional scene depiction. This method utilizes structure from motion (SfM) [6] to initialize 3D Gaussians and optimizes a 3D Gaussian as a volumetric model. Subsequently, it projects this model onto 2D surfaces to facilitate rasterization. 3D-GS uses an $\alpha$-blender to merge pixel colours effectively. This technique results in high-fidelity outputs with detailed resolution, enabling rendering at real-time speeds.
@@
Large-scale Scene Reconstruction    Concurrently, as Neural Radiance Fields (NeRF) [4] gain prominence for generating photorealistic perspectives in contemporary visual synthesis, a plethora of adaptations have surfaced. These aim to increase reconstruction quality [34, 16, 35, 36, 37, 38, 39, 40], accelerate rendering [10, 14, 13, 17, 18], and extend capabilities to dynamic scenarios [41, 42, 43, 44]. Among these, several methods [45, 46, 47, 48] have scaled NeRF to accommodate expansive scenes. Specifically, Block-NeRF [45] segments urban landscapes into several blocks, assigning view-specific training based on geographic location. Alternatively, Mega-NeRF [46] introduces a grid-oriented partitioning technique, linking each image pixel to various grids intersected by its corresponding ray. Different from heuristic partitioning methods, Switch-NeRF [48] has pioneered a mixture-of-experts NeRF framework to master scene segmentation. Conversely, Grid-NeRF [47] synergizes NeRF-based and grid-based strategies without segmenting the scene. Despite these improvements significantly elevating rendering precision over conventional methods, they often render slowly and lack finer details. In a recent development, 3D Gaussian Splatting (3DGS) [5] has been introduced. It provides an explicit, high-definition 3D representation that supports real-time rendering. However, these traditional 3DGS methods [5, 7, 8, 9] have been shown to consume significant resources when applied to extensive scenes, such as urban environments or scenic landscapes. This is primarily due to the considerable memory and graphics memory demands necessary for initial scene processing and the creation of Gaussian spheres. Previous methodologies [5, 7, 8, 49] for reconstructing large scenes typically relied on selecting a subset of images for training and then regenerating point clouds and viewpoints using COLMAP [50], which employs Structure-from-Motion (SfM) and Multi-View Stereo (MVS) techniques to derive camera positions and 3D structures from images. However, this approach proved to be inherently non-generalizable. The primary issue was the lack of effective scene segmentation, which led to random retention of images. Consequently, this resulted in fragmented reconstruction outcomes. Moreover, these approaches lead to disparate Gaussian outcomes, which could not be merged effectively. Each batch of partial images remained with isolated training results that lacked collective significance. Additionally, the use of incomplete image sets in training often resulted in inadequate COLMAP [50] results due to the failure to accurately select all required viewpoints for a comprehensive scene reconstruction. Our GaussianFocus successfully overcomes the limitations of the 3DGS-based methods [5, 7, 8] in training large-scale scenes through the introduction of innovative designs that efficiently subdivide, optimize, and integrate these scenes.
@@ 3 Methodology
Figure 2: Overview of GaussianFocus: Our model will monitor the size of Gaussian spheres during initialization and training. Constraints are applied to the scaling matrix $S$ within the covariance matrix to prevent Gaussian spheres’ excessive growth. Subsequently, the rendered image is divided into 64 parts. Each part independently calculates its attention values, which are then concatenated to form a comprehensive attention map. This map is multiplied back onto the original rendered image to produce an attention-enhanced image. Finally, this enhanced image and the original rendered image undergo multiple loss calculations against the ground truth. These include reconstruction ($L_{1}$), structural similarity ($L_{D-SSIM}$), edge ($L_{Edge}$), and frequency ($L_{Frequency}$) losses.
@@
3.1 Preliminary
In the foundational aspects of the 3DGS framework [5], the scene is represented using anisotropic 3D Gaussians that integrate differential properties typical of a volume-based approach but are rendered more effectively through a grid-based rasterization technique. Beginning with a collection of structure-from-motion (SfM) [6] data points, each point is established as the centroid ($\mu$) for a 3D Gaussian. The formula for a 3D Gaussian $G(x)$ is given by:
$G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\right),$ (1)
where $x$ represents a point within the 3D space, and $\Sigma$ represents the Gaussian covariance matrix which is constructed using
$\Sigma=RSS^{T}R^{T}.$ (2)
This configuration is derived from a scaling matrix $S$ and a rotational matrix $R$, guaranteeing its positivity and semi-definiteness.
Each Gaussian is not only linked with a colour $c_{i}$, defined through spherical harmonics but also paired with an opacity $\alpha$, impacting the merging process in rendering. Diverging from classic volumetric methods that employ ray-marching, this model projects 3D Gaussians onto a 2D plane $G^{2D}(x)$ and processes them through a grid-based rasterizer for sorting and $\alpha$-blending. The $\alpha$-blending formula is specified as
$C(x^{\prime})=\sum_{i\in K}c_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{j}),$ (3)
where $\sigma_{i}=\alpha_{i}G^{2D}_{i}(x^{\prime})$, $x^{\prime}$ represents the specified pixel position, and $K$ counts the Gaussians for that specified pixel in two dimensions. This approach facilitates the direct learning and optimization of the Gaussian features through a trainable differentiable rasterizer.
@@ 3.1 Preliminary
Given the significant computational demands, it is impractical to directly compute attention values for the entire rendered image due to the extensive data processing involved. Instead, both model-rendered image $P_{i}$ and the Ground Truth images $G_{i}$ are segmented into 8x8 regions to manage computational complexity effectively. For each segment of $P_{i}$, a query vector $q_{ij}$ is extracted using a 2D convolutional layer which is designed to capture detailed features and spatial relationships within the segment. Correspondingly, the key $k_{ij}$ and value $v_{ij}$ for each segment $j$ of $G_{i}$ are derived through similar 2D convolutional layers. These steps ensure that the essential components for the multi-head attention mechanism—queries, keys, and values (QKV)—are accurately assembled based on localized image features. The attention weights $w_{ij}$ for each segment can be calculated using the following equation:
@@
where the sum over $j$ aggregates the contributions of all segments into a unified attention profile for the entire image. This comprehensive attention map $A_{i}$ is then used to produce an attention-enhanced image $P_{i}^{{}^{\prime}}$ by element-wise multiplying it with the rendered image $P_{i}$:
@@
which enhances the original image by amplifying features that are deemed significant based on the attention mechanism. To further enhance the reconstruction’s accuracy, we compute edge loss $L_{Edge}$ and frequency loss $L_{Frequency}$ for this enhanced image in conjunction with the ground truth image. These losses are calculated alongside the standard loss comparisons between the original rendered image and the ground truth image. They will be discussed in Section 3.5.
@@ 3.3 Gaussian Sphere Constraints
During the initialization of Gaussian spheres, we impose constraints on the scaling matrix $S$ to control the covariance matrix’s influence, essential for accurately modelling spatial relationships in the scene. The adjustment of $S$ is dictated by the density of the initial point cloud data: for denser point clouds, we set a lower initial scaling value to reduce overlaps and redundancy, while for sparser distributions, we increase it to ensure sufficient scene coverage. This careful calibration of scaling factors helps maintain an optimal balance between preserving detail and enhancing computational efficiency. The scaling matrix constraint is defined as follows:
$S_{i}=S_{i}\cdot\alpha,\quad\text{if }S_{i}>\tau,\$ (8)
@@
Selective Split($\text{Selective Split}\;(S_{\gamma}),\quad\text{if }S_{\gamma}>\Omega$ (9)
where $S_{\gamma}$ denotes the scales of the scaling matrix of aged Gaussians and $\Omega$ is the threshold set to identify old Gaussians that require reevaluation. We set $\Omega$ = 0.3 in our experiment. These strategies ensure that our method maintains a balanced approach to managing the size and number of Gaussians within the 3DGS framework.
@@ 3.4 Subdivision-Based Reconstruction of Large Scenes
In response to 3DGS challenges [5, 7, 8], our method initiates a preprocessing step to acquire initial points from Structure-from-Motion (SfM) [6] of the large scene. As shown in Fig. 3, a three-dimensional bounding box is then constructed to encompass all initial point clouds. We divide this bounding box along its $xyz$ axes into $n\times n\times n$ blocks, where each block is defined to contain its respective subset of point clouds:
@@
Iterate $\text{Iterate }\forall pc\in\text{Point Cloud}:\text{if }\text{dist}(pc,\text{Block}_{ijk})>\theta\text{ then discard }pc,$ (11)
where $\text{dist}(\cdot)$ calculates the distance from the point to the nearest block boundary, and $\theta$ is a threshold value defining the maximum allowable distance for inclusion. Corresponding camera and Structure-from-Motion (SfM) points associated with each block are classified to assemble the essential initial files required for training. Each block undergoes independent training. The process concludes with the recombination of the divided scene’s 3D files, thus completing the reconstruction of the entire large scene. This modular approach alleviates the computational and memory constraints typically linked with large-scale scene reconstruction. By employing this method, we efficiently manage large scene datasets and enhance the scalability of our reconstruction processes.

### v2

@@ Abstract
Recent developments in 3D reconstruction and neural rendering have significantly propelled the capabilities of photo-realistic 3D scene rendering across various academic and industrial fields. The 3D Gaussian Splatting technique, alongside its derivatives, integrates the advantages of primitive-based and volumetric representations to deliver top-tier rendering quality and efficiency. Despite these advancements, the method tends to generate excessive redundant noisy Gaussians overfitted to every training view, which degrades the rendering quality. Additionally, while 3D Gaussian Splatting excels in small-scale and object-centric scenes, its application to larger scenes is hindered by constraints such as limited video memory, excessive optimization duration, and variable appearance across views. To address these challenges, we introduce GaussianFocus, an innovative approach that incorporates a patch attention algorithm to refine rendering quality and implements a Gaussian constraints strategy to minimize redundancy. Moreover, we propose a subdivision reconstruction strategy for large-scale scenes, dividing them into smaller, manageable blocks for individual training. Our results indicate that GaussianFocus significantly reduces unnecessary Gaussians and enhances rendering quality, surpassing existing State-of-The-Art (SoTA) methods. Furthermore, we demonstrate the capability of our approach to effectively manage and render large scenes, such as urban environments, whilst maintaining high fidelity in the visual output. Please visit https://github.com/HZXu-526/GaussianFocus for code.
@@ 1 Introduction
Novel View Synthesis (NVS) is fundamental for modern computer graphics and vision, extending to virtual reality, autonomous driving, and robotics. Primitive-based models such as meshes and point clouds [12, 17, 28], optimized for GPU rasterization, deliver fast but often lower-quality images with discontinuities. The introduction of Neural Radiance Fields (NeRF) [15] marked a significant advancement, employing a multi-layer perceptron (MLP) to achieve high-quality, geometrically consistent renderings of novel viewpoints. However, NeRF’s reliance on time-consuming stochastic sampling can lead to reduced performance and potential noise issues.
Recent advancements in 3D Gaussian Splatting (3DGS) [9] have significantly enhanced rendering quality and speed. This technique refines a set of 3D Gaussians, initialized using Structure from Motion (SfM) [20], to model scenes with inherent volumetric continuity. This facilitates fast rasterization by projecting them onto 2D planes. However, 3DGS often produces artifacts when camera viewpoints deviate from the training set and lack detail during zooming. To address these issues, newer models [30, 14] employ a 3D smoothing filter to regularize the frequency distribution and utilize anchor points to initialize 3D Gaussians, thereby enhancing visual accuracy and applicability in diverse scenarios. Despite these advances, 3DGS-based models still tend to produce oversized Gaussian spheres that ignore scene structure, leading to redundancy and scalability issues in complex environments. Additionally, these models struggle with detail reconstruction, particularly at edges and high-frequency areas. This often leads to suboptimal rendering quality. Moreover, reconstructing large-scale scenes like towns or cities represents a significant challenge due to GPU memory constraints and computational demands. To mitigate these problems, models often reduce the number of training inputs randomly, which compromises reconstruction quality and results in incomplete outcomes.
@@
    We propose a 3DGS-based patch attention algorithm with novel edge and frequency losses designed to enhance details and reduce spatial frequency artifacts within scene reconstructions. This improves the detailing quality and intricacy of the rendered scenes.
@@
    We impose constraints on overly large Gaussian spheres to mitigate the occurrence of “air walls”, thus improving the scene reconstruction fidelity and enhancing the granularity of the resulting models. Moreover, these constraints allow the achievement of superior reconstruction results with fewer training iterations.
@@
    For large-scale scene reconstruction, our approach involves subdividing the scene for reconstruction and subsequent recombination. This method addresses the challenge posed by existing 3DGS-based models that fail to directly reconstruct extensive scenes, thereby enhancing the scalability and applicability of our reconstruction framework.
@@ 2 Related Work
Volumetric Rendering methods Volumetric approaches represent scenes as continuous functions specifying volume, density, and color characteristics. The introduction of Neural Radiance Fields (NeRF) [15] was a pivotal advancement, using a multilayer perceptron (MLP) to parameterize these functions and create photorealistic images that vary with viewer perspective through volumetric ray tracing. However, the computational and memory demands of vanilla NeRF limit its practical application. Subsequent research has improved NeRF’s efficiency and scalability through discretized or sparse volumetric frameworks like voxel grids and hash tables [4, 21, 16], which hold learnable features acting as positional encodings for 3D coordinates. Additionally, some methods employ hierarchical sampling techniques [2, 18, 29] and utilize low-rank approximations [4]. Despite these enhancements, the dependence on volumetric ray marching continues, which leads to compatibility challenges with traditional graphics equipment and systems primarily engineered for polygonal rendering.
@@
Large-scale Scene Reconstruction Concurrently, a plethora of adaptations have surfaced as Neural Radiance Fields (NeRF) [15] gain prominence for generating photorealistic perspectives in contemporary visual synthesis. These aim to increase reconstruction quality [1, 2, 3, 24, 25], accelerate rendering [4, 16], and extend capabilities to dynamic scenarios [27, 6]. Among these, several methods [22, 23] have scaled NeRF to accommodate expansive scenes. However, they often render slowly and lack finer details. The newly introduced 3D Gaussian Splatting (3DGS) [9] enables explicit, high-definition 3D representations with real-time rendering capabilities. However, applying traditional 3DGS methods [9, 30, 14, 7] to large scenes, such as urban or scenic landscapes, demands significant memory and graphics resources for initial scene processing and Gaussian sphere creation. Several works [13, 10] have scaled 3DGS to large-scale scenes, but their methods fail to adapt to large changes in viewpoint. Previously, methods [9, 30, 14] could not directly process large-scale scenes and instead relied on selecting a reduced subset of input images to perform pre-processing using COLMAP [19], which applies Structure-from-Motion (SfM) and Multi-View Stereo (MVS) to estimate camera poses and sparse point clouds. However, this strategy often led to incomplete or fragmented scene reconstructions due to suboptimal scene coverage and loss of important visual information. Moreover, using incomplete image sets often resulted in poor COLMAP [19] outcomes, failing to cover all necessary viewpoints for comprehensive scene reconstruction. Our GaussianFocus successfully overcomes the limitations of the 3DGS-based methods [9, 30, 14] in training large-scale scenes through the introduction of methods that efficiently subdivide, optimize, and integrate these scenes.
@@ 3 Methodology
Figure 2: Overview of GaussianFocus: Our model monitors the size of Gaussian spheres during initialization and training. Constraints are applied to the scaling matrix $S$ within the covariance matrix to prevent the excessive growth of the Gaussian spheres. Subsequently, the rendered image is divided into 64 parts. Each part independently calculates its attention values, which are then concatenated to form a comprehensive attention map. This map is multiplied back onto the original rendered image to produce an attention-enhanced image. Finally, this enhanced image and the original rendered image undergo multiple loss calculations against the ground truth. These include reconstruction ($L_{1}$), structural similarity ($L_{D-SSIM}$), edge ($L_{Edge}$), and frequency ($L_{Frequency}$) losses.
@@
Given the significant computational demands, it is impractical to directly compute attention values for the entire rendered image due to the extensive data processing involved. Instead, both model-rendered image $P_{i}$ and the Ground Truth images $G_{i}$ are segmented into 8x8 regions to manage computational complexity effectively. For each segment of $P_{i}$, a query vector $q_{ij}$ is extracted using a 2D convolutional layer, which is designed to capture detailed features and spatial relationships within the segment. Correspondingly, the key $k_{ij}$ and value $v_{ij}$ for each segment $j$ of $G_{i}$ are derived through similar 2D convolutional layers. These steps ensure that the essential components for the multi-head attention mechanism—queries, keys, and values (QKV)—are accurately assembled based on localized image features. The attention weights $w_{ij}$ for each segment can be calculated using the following equation:
@@
where the sum over $j$ aggregates and assembles the contributions of all segments into a unified attention map for the entire image. This comprehensive attention map $A_{i}$ is then used to produce an attention-enhanced image $P_{i}^{{}^{\prime}}$ by element-wise multiplying it with the rendered image $P_{i}$:
@@
which enhances the original image by amplifying features that are deemed significant based on the attention mechanism. To further enhance the reconstruction accuracy, we compute edge loss $L_{Edge}$ and frequency loss $L_{Frequency}$ for this enhanced image in conjunction with the ground truth image. These losses are calculated alongside the standard loss comparisons between the original rendered image and the ground truth image. They will be discussed in Section 3.4.
@@ 3.2 Gaussian Sphere Constraints
During the initialization of Gaussian spheres, we impose constraints on the scaling matrix $S$ to control the covariance matrix’s influence, essential for accurately modelling spatial relationships in the scene. The adjustment of $S$ is dictated by the density of the initial point cloud data: for denser point clouds, we set a lower initial scaling value to reduce overlaps and redundancy, while for sparser distributions, we increase it to ensure sufficient scene coverage. We used the distance to determine the density for the point clouds. This careful calibration of scaling factors helps maintain an optimal balance between preserving detail and enhancing computational efficiency. The scaling matrix constraint is defined as follows:
$S_{i}=S_{i}\cdot\alpha,\quad\text{if }S_{i}>\tau,\\$ (5)
@@
$\text{Selective Split}\;(S_{\gamma}),\quad\text{if }S_{\gamma}>\Omega$ (6)
where $S_{\gamma}$ denotes the scales of the scaling matrix of aged Gaussians and $\Omega$ is the scale threshold set to identify old Gaussians that require reevaluation. We set $\Omega$ = 0.3 in our experiment. This strategy reduces the unnecessary split in training. These strategies ensure that our method maintains a balanced approach to managing the size and number of Gaussians within the 3DGS framework.
@@ 3.3 Subdivision-Based Reconstruction of Large Scenes
In response to 3DGS challenges [9, 30, 14], our method initiates a preprocessing step to acquire initial points from Structure-from-Motion (SfM) [20] of the large scene. As shown in Fig. 3, a three-dimensional bounding box is then constructed to encompass initial point clouds. We divide this bounding box along its xyzxyz axes into $n\times n\times n$ blocks, where each block is defined to contain its respective subset of point clouds:
@@
$\text{Iterate }\forall pc\in\text{Point Cloud}:\text{if }\text{dist}(pc,\text{Center}_{ijk})>\theta\text{ then discard }pc,$ (8)
where $\text{dist}(\cdot)$ calculates the distance from the point to the block center, and $\theta$ is a threshold value defining the maximum allowable distance for inclusion. Corresponding camera and Structure-from-Motion (SfM) points associated with each block are classified to assemble the essential initial inputs required for training. Each block undergoes independent training. The process concludes with the recombination of the divided scene’s 3D files, thus completing the reconstruction of the entire large scene. This modular approach alleviates the computational and memory constraints typically linked with large-scale scene reconstruction. By employing this method, we efficiently manage large scene datasets and enhance the scalability of our reconstruction processes.